You have 8 balls. Each ball is labeled 1,2,3,4,5,6,7,8 and are inside of one bag. Imagine you randomly take 3 balls from the bag. What is the probability that the product of the three numbers on the balls you select is even?
Answer #1

The balls are numbered 1,2,3,4,5,6,7,8.

3 balls are chosen at random.

4 of them are odd and 4 are even.

The product will be even if minimum one of the balls numbered is even.

Total number of ways 3 balls can be drawn from a bag of 8 balls without replacement = 8P3 = 8! /(8-3)! = 8! / 5! = 8*7*6 = 336

Total number of ways the product is odd = Total ways all three balls drawn are odd = 4P3 = 4

Total number of ways the product is even = Total number of ways 3 balls can be drawn from a bag of 8 balls without replacement - Total number of ways the product is odd

Total number of ways the product is even = 336 - 4 = 332

P[ the product is even ] = Total number of ways the product is even / Total number of ways 3 balls can be drawn from a bag of 8 balls

P[ the product is even ] = 332/336 = 0.98809 = 0.99

P[ the product is even ] = 0.99